In a lottery game, a single ball is drawn at random from a container that contains 25 identical balls numbered from 1 – 25. Compute the probability that the number drawn is less than 6 or greater than 19.
Answer by stanbon(75887)

In a lottery game, a single ball is drawn at random from a container that contains 25 identical balls numbered from 1 – 25. Compute the probability that the number drawn is less than 6 or greater than 19.
----
# of ways to succeed:: 5+6 = 11
-----
P(<6 OR >19) = P(<6)+P(>19) = (5+6)/25 = 11/25
--------------
Cheers,
Stan H.
--------------